# Probing Neural Topology of Large Language Models

### v2

@@ Abstract
Probing large language models (LLMs) has yielded valuable insights into their internal mechanisms by linking neural activations to interpretable semantics. However, the complex mechanisms that link neuron’s functional co-activation with the emergent model capabilities remains largely unknown, hindering a deeper understanding and safer development of LLMs. In this work, we introduce graph probing, a method for uncovering the functional connectivity of LLM neurons and relating it to language generation performance. By probing models across diverse LLM families and scales, we discover a universal predictability of next-token prediction performance using only neural topology, which persists even when retaining just 1% of neuron connections. Strikingly, probing on topology outperforms probing on activation by up to 130.4%, suggesting that neural topology contains orders of richer information of LLM performance than neural activation, which can be easily extracted with simple linear or MLP probes. To explain the dependence between neural topology and language performance, we identify default networks and hub neurons in LLMs and provide causal evidence by interventional experiments on multiple benchmarks, showing that LLMs actually exploit these topological information. Further analyses suggest that neural topology can be effectively leveraged to improve the efficiency, reliability, and safety of LLMs through proof-of-concept applications in model pruning, hallucination detection, and LLM fingerprinting. Codes and data for the graph probing toolbox are available at https://github.com/DavyMorgan/llm-graph-probing.
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable generative capabilities (Wei et al., 2022; Touvron et al., 2023; GLM et al., 2024; Team et al., 2024; Guo et al., 2025), yet our understanding of how they succeed and what they have learned remains limited (Sharkey et al., 2025). Probing, which extract interpretable features from neural activations (Alain & Bengio, 2017), has emerged as a powerful approach for reverse-engineering LLMs (Belinkov, 2022; Gurnee & Tegmark, 2024). For instance, Gurnee et al.(Gurnee & Tegmark, 2024) showed that LLMs encode a compact world model of space and time using linear regression probes. Unsupervised probing, such as sparse auto-encoders (Engels et al., 2025; Gao et al., 2025; Rajamanoharan et al., 2024; Lieberum et al., 2024; Mudide et al., 2025) and cross-layer transcoders (Dunefsky et al., 2024; Lindsey et al., 2025), have further revealed dictionaries of interpretable, mono-semantic concepts (Huben et al., 2024) and even causal circuits (Marks et al., 2025), corresponding to directions in neural latent space. While these advances shed light on the semantics of neural activations (Sharkey et al., 2025), much less is known about how neurons are functionally connected, i.e. the neural topology, which is believed to play an essential role in the emergence of intelligence (Rathi et al., 2025; Bassett & Sporns, 2017).
@@
Recent studies have drawn compelling parallels between neurons in LLMs and those in the human brain (Toneva & Wehbe, 2019; Schrimpf et al., 2021; Caucheteux et al., 2023; Kumar et al., 2024; Rathi et al., 2025; Mischler et al., 2024; Tuckute et al., 2024; Bonnasse-Gahot & Pallier, 2024; Sun et al., 2024; Liu et al., 2025), revealing shared properties such as spatial-functional organization (Kumar et al., 2024; Rathi et al., 2025) and left lateralization (Bonnasse-Gahot & Pallier, 2024). Neural activations at internal layers of LLMs have also been shown to reliably predict human brain responses given the same linguistic stimuli (Schrimpf et al., 2021; Tuckute et al., 2024; Luo et al., 2022). However, these efforts primarily focus on static neural activations of LLMs, while overlooking the key aspect of temporal and functional neural topology that has been studied in neuroscience for decades (Bassett & Bullmore, 2006; Bassett & Sporns, 2017; Fotiadis et al., 2024). Moreover, although analogies between LLMs and human brains are insightful (Toneva & Wehbe, 2019; Goldstein et al., 2022), few works explicitly connect these findings to the language generation performance, which is one of the primary indicators of an LLM’s intelligence.
@@
In this work, we introduce graph probing, a novel approach for investigating the functional connectivity of neurons in LLMs and its relationship to language generation performance. By analyzing neural activity time series as LLMs process text token by token, we compute temporal and functional correlations between neurons to construct dynamic neural graphs. Using this large-scale dataset of text-induced neural topology, we train probes to predict LLMs’ accuracy in auto-regressively generating the corresponding text. In essence, graph probing connects the micro-level topology of how neurons are connected given a token sequence, to the macro-level performance of how well LLMs predict these tokens, offering a new lens to study the emergent capabilities of LLMs. Our method is summarized in Figure 1 and described in Section 2.
@@
We then apply our graph probing framework to comprehensively analyze the neural topology of LLMs through extensive experiments. First, we demonstrate that auto-regressive language generation performance can be reliably predicted using only the neural connectivity graph. This predictability holds universally across LLM families and scales, outperforming activation-based probing approaches by up to 130.4% even when preserving only 1% of neuron connections, with empirical results spanning GPT (Radford et al., 2019), Pythia (Biderman et al., 2023), and Qwen (Yang et al., 2024), ranging from millions to billions of parameters (Section 3). Next, we show causal evidence through interventional experiments and analysis on the MMLU (Hendrycks et al., 2021; Wang et al., 2024) benchmark, discovering stable default neural topology and hub neurons in LLMs, and more importantly, validating that LLMs actually utilize their internal topological information when generating text (Section 4). Finally, we offer three proof-of-concept applications of neural topology, showcasing its potential in model pruning, hallucination detection, and LLM fingerprinting (Section 5). While not without limitations, we expect graph probing to provide valuable insights into the inner workings of LLMs and to guide their future development in an interpretable and safe way.
@@ 2 Graph Probing
To construct neural graphs from LLMs, we draw inspiration from neuroscience where functional brain networks are derived from temporal correlations in fMRI or EEG activation signals (Bassett & Sporns, 2017; Vértes et al., 2012; Bullmore & Bassett, 2011), as shown in Figure 1. Given an LLM composed of stacked attention layers, the neural topology is constructed as follows:
Neural Activity: $\mathbf{H}$ $=\mathtt{HIDDEN\_STATE}(\mathtt{LLM}(X))=[\mathbf{h_{0}},\mathbf{h_{1}},\ldots,\mathbf{h_{t}}]\in\mathbb{R}^{n\times t},$ (1)
Neural Topology: $\mathbf{A}$ $=\bigl(a_{ij}\bigr)\in\mathbb{R}^{n\times n},$ (2)
$a_{ij}$ $=\rho(\mathbf{H}_{i,:},\mathbf{H}_{j,:})=\frac{\sum_{k=0}^{t}{(\mathbf{H}_{i,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{i,:}}\mskip-1.5mu}\mskip 1.5mu)(\mathbf{H}_{j,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{j,:}}\mskip-1.5mu}\mskip 1.5mu)}}{\sqrt{\sum_{k=0}^{t}{(\mathbf{H}_{i,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{i,:}}\mskip-1.5mu}\mskip 1.5mu)^{2}}}\sqrt{\sum_{k=0}^{t}{(\mathbf{H}_{j,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{j,:}}\mskip-1.5mu}\mskip 1.5mu)^{2}}}},$ (3)
where neurons at each layer produce a time series of hidden states $\mathbf{H}$ as the model processes a token sequence $X=[x_{0},x_{1},\ldots,x_{t}]$, and the temporal co-activation patterns among neurons define their functional connectivity. We capture this through a complete $n\times n$ weighted connectivity matrix $\mathbf{A}$, where each node corresponds to a neuron and each edge weight $a_{ij}$ represents the Pearson correlation coefficient between the activation time series of a pair of neurons $i$ and $j$ (Bassett & Sporns, 2017; Fotiadis et al., 2024).
@@
We propose graph probing to study the dependence between LLM performance and neural topology. Specifically, we adopt simple linear or multi-layer perceptrons (MLP) probes (Rumelhart et al., 1986) that take neural topology as input to predict its corresponding language generation performance, as illustrated in Figure 1. Given a connectivity matrix $\mathbf{A}$ induced by feeding a tokenized sequence $X$ to an LLM, where each element $a_{ij}$ denotes the functional connectivity (Pearson correlation coefficient) between neurons $i$ and $j$, our probe produces the graph representation $\mathbf{z}$ as follows:
@@
where $\mathtt{Flatten}(\mathbf{A})\in\mathbb{R}^{(n\times n)}$ is the flattened topology matrix, $\mathbf{W_{1}}\in\mathbb{R}^{1\times(n\times n)},\mathbf{W_{2}}\in\mathbb{R}^{d\times(n\times n)},\mathbf{W_{3}}\in\mathbb{R}^{1\times d}$ are weight matrices ($d$ as a hyper-parameter), $\hat{y}$ is the prediction of language performance. In essence, the core and only difference of our method lies at the input, where we probe on neural topology instead of neural activation by existing approaches (Belinkov, 2022; Gurnee et al., 2023). We will later show that neural topology contains orders of richer information than neural activation regarding LLMs’ intelligence.
@@
By its definition, graph probing can be utilized to predict almost any fact of interest, and in this work we test it on various performance-related metrics, including perplexity, hallucination, and functional specialization. Here we introduce the case of perplexity as it is a fundamental metric directly reflecting LLMs’ language generation performance, while details of other cases are provided in later sections. Specifically, given a token sequence, the LLM’s auto-regressive generation performance is commonly measured by perplexity (Bengio et al., 2003) which corresponds to the exponentiated average negative log-likelihood over the token sequence:
$\textbf{Perplexity:}\quad\mathtt{PPL}(X)=\exp\left\{-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})\right\}.$ (6)
@@ 3 Results
Evaluation. We split the dataset into training and test sets using an 8:2 ratio. Having learned graph probes on the training set, we evaluate their out-of-sample prediction performance on the test set, which reveals the extent to which micro-level neural topology is predictive of macro-level language generation ability. To quantify the effectiveness of graph probing, we report standard regression metrics on our test data, including mean squared error (MSE), mean absolute error (MAE), coefficient of determination ($R^{2}$), Pearson correlation ($\rho_{p}$), and Spearman rank correlation ($\rho_{s}$). We compare with existing activation based baselines using both linear and MLP probes (Alain & Bengio, 2017; Belinkov, 2022; Gurnee & Tegmark, 2024; Tuckute et al., 2024).
@@
Sparsity and scalability. Probing on complete graphs, i.e., dense $n\times n$ connectivity matrices that capture pairwise functional correlations between all neurons, can become computationally prohibitive as the LLM size increases, due to the quadratic number of edges that directly impacts the computational cost in both time and memory. For instance, while complete graph probing is feasible for Pythia-160M with 768 neurons and 0.6M edges per layer, the number of edges in Qwen2.5-14B–comprising 5,120 neurons per layer–explodes to over 26M per graph. To address this, on the one hand, we investigate whether perplexity can still be predicted from sparse graphs with weakly correlated edges pruned out by thresholding, which is commonly employed in human brain network construction (Bassett & Sporns, 2017); on the other hand, we develop graph neural network probe (Kipf & Welling, 2017) that reduces the number of probe parameters from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\cdot d)$ by weight sharing (details in Appendix A.4). We train graph probes on neural topology with varying levels of sparsification (Figure 3(a)). Surprisingly, the predictive performance remains remarkably stable even after removing up to 90% of the edges, with minimal degradation. Notably, even under extreme sparsity where only 1% of the original edges are retained, the neural topology still enables effective prediction of perplexity, achieving above 0.6 $R^{2}$ that is still higher than activation probing.
@@
The above experiments suggest that most of the predictive signal resides in a small subset of strong connections, making it possible to significantly reduce the number of edges while preserving nearly all critical topological information. Leveraging this insight, we scale up graph probing to much larger models by operating on sparsified neural topology. While our earlier results focused on models with fewer than 0.5B parameters, we now train probes on sparse graphs derived from LLMs with up to 14B parameters. As shown in Figure 3(b), graph probing continues to exhibit strong regression performance across all six large models, achieving a maximum $R^{2}$ of over 0.91, providing compelling evidence that the relationship between neural topology and language modeling performance is universal across model sizes. Particularly, the gap between baselines and graph probing remains as substantial as 92.6%, confirming the informativeness of neural topology.
@@
Probing different layers. We further train probes on neural topology derived from different layers of GPT-2, Pythia-160M, and Qwen2.5-0.5B models. Figure 3(c) reveals that while topologies from all layers are predictive of LLM performance, the middle layers are consistently the most informative across all three models, with $R^{2}$ values exceeding 0.8. This finding aligns with prior work identifying the middle layers of LLMs as the hub of semantic processing, whereas the initial and final layers are more specialized for {de, re}-tokenization (Gurnee & Tegmark, 2024; Tuckute et al., 2024).

### v3

@@ Abstract
Probing large language models (LLMs) has yielded valuable insights into their internal mechanisms by linking neural activations to interpretable semantics. However, the complex mechanisms that link neuron’s functional co-activation with the emergent model capabilities remains largely unknown, hindering a deeper understanding and safer development of LLMs. In this work, we introduce graph probing, a method for uncovering the functional connectivity of LLM neurons and relating it to language generation performance. By probing models across diverse LLM families and scales, we discover a universal predictability of language generation and understanding performance using only neural topology, which persists even when retaining just 1% of neuron connections. Strikingly, probing on topology outperforms probing on activation by up to 130.4% and 67.7% on perplexity and space/time semantic regression respectively, suggesting that neural topology contains orders of richer information of LLM performance than neural activation, which can be easily extracted with simple linear or MLP probes. To explain the dependence between neural topology and language performance, we identify default networks and hub neurons in LLMs and provide causal evidence by interventional experiments on multiple benchmarks, showing that LLMs actually exploit these topological information. Further analyses suggest that graph probing can be effectively leveraged to improve the efficiency and reliability of LLMs through proof-of-concept applications in model pruning and hallucination detection. Codes and data for the graph probing toolbox are available at https://github.com/DavyMorgan/llm-graph-probing.
Large Language Models, Interpretability, Neural Topology, Probing
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable generative capabilities (Wei et al., 2022; Touvron et al., 2023; GLM et al., 2024; Team et al., 2024; Guo et al., 2025), yet our understanding of how they succeed and what they have learned remains limited (Sharkey et al., 2025). Probing, which extract interpretable features from neural activations (Alain and Bengio, 2017; Rogers et al., 2020; Hewitt and Liang, 2019; Voita and Titov, 2020; Pimentel et al., 2020), has emerged as a powerful approach for reverse-engineering LLMs (Belinkov, 2022; Gurnee and Tegmark, 2024). For instance, Gurnee et al.(Gurnee and Tegmark, 2024) showed that LLMs encode a compact world model of space and time using linear regression probes. Unsupervised probing, such as sparse auto-encoders (Engels et al., 2025; Gao et al., 2025; Rajamanoharan et al., 2024; Lieberum et al., 2024; Mudide et al., 2025) and cross-layer transcoders (Dunefsky et al., 2024; Lindsey et al., 2025), have further revealed dictionaries of interpretable, mono-semantic concepts (Huben et al., 2024) and even causal circuits (Marks et al., 2025), corresponding to directions in neural latent space. While these advances shed light on the semantics of neural activations (Sharkey et al., 2025), much less is known about how neurons are functionally connected, i.e. the neural topology, which is believed to play an essential role in the emergence of intelligence (Rathi et al., 2025; Bassett and Sporns, 2017).
@@
Recent studies have drawn compelling parallels between neurons in LLMs and those in the human brain (Toneva and Wehbe, 2019; Schrimpf et al., 2021; Caucheteux et al., 2023; Kumar et al., 2024; Rathi et al., 2025; Mischler et al., 2024; Tuckute et al., 2024; Bonnasse-Gahot and Pallier, 2024; Sun et al., 2024a; Liu et al., 2025), revealing shared properties such as spatial-functional organization (Kumar et al., 2024; Rathi et al., 2025) and left lateralization (Bonnasse-Gahot and Pallier, 2024). Neural activations at internal layers of LLMs have also been shown to reliably predict human brain responses given the same linguistic stimuli (Schrimpf et al., 2021; Tuckute et al., 2024; Luo et al., 2022). However, these efforts primarily focus on static neural activations of LLMs, while overlooking the key aspect of temporal and functional neural topology that has been studied in neuroscience for decades (Bassett and Bullmore, 2006; Bassett and Sporns, 2017; Fotiadis et al., 2024). Moreover, although analogies between LLMs and human brains are insightful (Toneva and Wehbe, 2019; Goldstein et al., 2022), few works explicitly connect these findings to the language generation performance, which is one of the primary indicators of an LLM’s intelligence.
@@
In this work, we introduce graph probing, a novel approach for investigating the functional connectivity of neurons in LLMs and its relationship to language generation and understanding performance. By analyzing neural activity time series as LLMs process text token by token, we compute temporal and functional correlations between neurons to construct dynamic neural graphs. Using this large-scale dataset of text-induced neural topology, we train probes to predict LLMs’ accuracy in auto-regressively generating the corresponding text, as well as how LLMs represent fundamental concepts like space and time. In essence, graph probing connects the micro-level topology of how neurons are connected given a token sequence, to the macro-level performance of how well LLMs understand and predict these tokens, offering a new lens to study the emergent capabilities of LLMs. Our method is summarized in Figure 1 and described in Section 2.
@@
We then apply our graph probing framework to comprehensively analyze the neural topology of LLMs through extensive experiments. First, we demonstrate that language understanding and generation performance can be reliably predicted using only the neural connectivity graph. This predictability holds universally across LLM families and scales, outperforming activation-based probing approaches by up to 130.4% even when preserving only 1% of neuron connections, with empirical results spanning GPT (Radford et al., 2019), Pythia (Biderman et al., 2023), and Qwen (Yang et al., 2024), ranging from millions to billions of parameters (Section 3). Next, we show causal evidence through interventional experiments and analysis on the MMLU (Hendrycks et al., 2021; Wang et al., 2024) benchmark, discovering stable default neural topology and hub neurons in LLMs, and more importantly, validating that LLMs actually utilize their internal topological information when generating text (Section 4). Finally, we offer two proof-of-concept applications of graph probing, showcasing its potential in model pruning and hallucination detection (Section 5). While not without limitations, we expect graph probing to provide valuable insights into the inner workings of LLMs and to guide their future development in an interpretable and reliable way.
@@ 2 Graph Probing
To construct neural graphs from LLMs, we draw inspiration from neuroscience where functional brain networks are derived from temporal correlations in fMRI or EEG activation signals (Bassett and Sporns, 2017; Vértes et al., 2012; Bullmore and Bassett, 2011), as shown in Figure 1. Given an LLM composed of stacked attention layers, the neural topology is constructed as follows:
$\mathbf{H}=[\mathbf{h_{0}},\mathbf{h_{1}},\ldots,\mathbf{h_{t}}]\in\mathbb{R}^{n\times t},$ (1)
$\mathbf{A}=\bigl(a_{ij}\bigr)\in\mathbb{R}^{n\times n},$ (2)
$a_{ij}=\rho(\mathbf{H}_{i,:},\mathbf{H}_{j,:})$ (3)
$=\frac{\sum_{k=0}^{t}{(\mathbf{H}_{i,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{i,:}}\mskip-1.5mu}\mskip 1.5mu)(\mathbf{H}_{j,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{j,:}}\mskip-1.5mu}\mskip 1.5mu)}}{\sqrt{\sum_{k=0}^{t}{(\mathbf{H}_{i,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{i,:}}\mskip-1.5mu}\mskip 1.5mu)^{2}}}\sqrt{\sum_{k=0}^{t}{(\mathbf{H}_{j,k}-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{H}_{j,:}}\mskip-1.5mu}\mskip 1.5mu)^{2}}}},$ (4)
where neurons at each layer produce a time series of hidden states $\mathbf{H}$ as the model processes a token sequence $X=[x_{0},x_{1},\ldots,x_{t}]$ ($n$ and $t$ represent the number of neurons and tokens), and the temporal co-activation patterns among neurons define their functional connectivity. We capture this through a complete $n\times n$ weighted connectivity matrix $\mathbf{A}$, where each node corresponds to a neuron and each edge weight $a_{ij}$ represents the Pearson correlation coefficient between the activation time series of a pair of neurons $i$ and $j$ (Bassett and Sporns, 2017; Fotiadis et al., 2024).
@@
We propose graph probing to study the dependence between LLM performance and neural topology. Specifically, we adopt simple linear or multi-layer perceptrons (MLP) probes (Rumelhart et al., 1986) that take neural topology as input to predict its corresponding language understanding performance, as illustrated in Figure 1. Given a connectivity matrix $\mathbf{A}$ induced by feeding a tokenized sequence $X$ to an LLM, where each element $a_{ij}$ denotes the functional connectivity (Pearson correlation coefficient) between neurons $i$ and $j$, our probe produces the graph representation $\mathbf{z}$ as follows:
@@
where $\mathtt{Flatten}(\mathbf{A})\in\mathbb{R}^{(n\times n)}$ is the flattened topology matrix, $\mathbf{W_{1}}\in\mathbb{R}^{1\times(n\times n)},\mathbf{W_{2}}\in\mathbb{R}^{d\times(n\times n)},\mathbf{W_{3}}\in\mathbb{R}^{1\times d}$ are weight matrices ($d$ as a hyper-parameter), $\hat{y}$ is the prediction of language performance. In essence, the core and only difference of our method lies at the input, where we probe on neural topology instead of neural activation by existing approaches (Belinkov, 2022; Gurnee et al., 2023). We will later show that topology contains orders of richer information than activation regarding LLMs’ intelligence.
@@
By its definition, graph probing can be utilized to predict almost any fact of interest, and in this work we test it on various performance-related metrics, including perplexity, hallucination, and functional specialization, as well as fundamental semantics like space and time. Here we introduce the case of perplexity (PPL) as it is a fundamental metric directly reflecting LLMs’ auto-regressive language generation performance (Bengio et al., 2003), while details of other cases are provided in later sections. Specifically, given a token sequence, the perplexity is commonly calculated as the exponentiated average negative log-likelihood over the token sequence:
$\mathtt{PPL}(X)=\exp\left\{-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})\right\}.$ (7)
@@ 3 Results
Evaluation. We split the dataset into training and test sets using an 8:2 ratio. Having learned graph probes on the training set, we evaluate their out-of-sample prediction performance on the test set, which reveals the extent to which micro-level neural topology is predictive of macro-level language generation ability. To quantify the effectiveness of graph probing, we report standard regression metrics on our test data, including mean squared error (MSE), mean absolute error (MAE), coefficient of determination ($R^{2}$), Pearson correlation ($\rho_{p}$), and Spearman rank correlation ($\rho_{s}$). We compare with activation based baselines using both linear and MLP probes (Alain and Bengio, 2017; Belinkov, 2022; Gurnee and Tegmark, 2024; Tuckute et al., 2024).
@@
Table 2: Probing results on Arts (time) and World Places (space) datasets. We compare Average Activation and our Graph Probing method ( indicates $p$-value << 0.05).
LLM Probe Arts (time) World Places (space)
MSE MAE $R^{2}$ MSE MAE $R^{2}$
GPT-2 Activation (Linear) 0.0406 0.1596 0.2945 0.0360 0.1957 0.5106
Activation (MLP) 0.0382 0.1511 0.3359 0.0311 0.1727 0.5733
Graph (Linear) 0.0326 0.1371 0.4341 0.0317 0.1765 0.5689
Graph (MLP) 0.0267 0.1230 0.5361 0.0258 0.1514 0.6486
Pythia-160M Activation (Linear) 0.0399 0.1580 0.3071 0.0360 0.1968 0.5116
Activation (MLP) 0.0391 0.1560 0.3211 0.0306 0.1705 0.5812
Graph (Linear) 0.0282 0.1275 0.5100 0.0267 0.1583 0.6335
Graph (MLP) 0.0266 0.1234 0.5387 0.0246 0.1447 0.6636
Qwen2.5-0.5B Activation (Linear) 0.0325 0.1406 0.4359 0.0296 0.1747 0.6002
Activation (MLP) 0.0314 0.1368 0.4555 0.0268 0.1574 0.6367
Graph (Linear) 0.0258 0.1221 0.5524 0.0215 0.1364 0.7080
Graph (MLP) 0.0268 0.1254 0.5341 0.0223 0.1398 0.6953
@@
Besides predicting perplexity, we also extend our analysis to include a wider range of semantics, probing neural topology to predict time and space context of the input text. We adopt the space and time datasets utilized by (Gurnee and Tegmark, 2024). We construct the neural topology for input texts regarding historical artworks and world places, and trained probes to predict: (1) time: the release year of historical artworks; (2) space: the longitude and latitude of world places. The results are summarized in Table 2, which serve two critical purposes. Firstly, they validate findings from previous work (Gurnee and Tegmark, 2024) showing that LLMs possess internal representations of time and space. Secondly and crucially, they confirm that neural topology contains significantly richer information than neural activation for these semantic tasks. The substantial performance margin of as much as 67.77% suggests that the structure of computation (how neurons connect) is more predictive than the magnitude of computation (activations) for extracting high-level concepts.
@@
Sparsity and scalability. Probing on complete graphs, i.e., dense $n\times n$ connectivity matrices that capture pairwise functional correlations between all neurons, can become computationally prohibitive as the LLM size increases, due to the quadratic number of edges that directly impacts the computational cost in both time and memory. For instance, while complete graph probing is feasible for Pythia-160M with 768 neurons and 0.6M edges per layer, the number of edges in Qwen2.5-14B–comprising 5,120 neurons per layer–explodes to over 26M per graph. To address this, on the one hand, we investigate whether probes can be applied to sparse graphs with weakly correlated edges pruned out by thresholding, which is commonly employed in human brain network construction (Bassett and Sporns, 2017) and such sparsity is frequently observed in artificial neural networks (Vig et al., 2020; Timkey and van Schijndel, 2021; Dettmers et al., 2022); on the other hand, we develop graph neural network probe (Kipf and Welling, 2017) that reduces the number of probe parameters from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\cdot d)$ by weight sharing (details in Appendix A.4). We train graph probes on neural topology with varying levels of sparsification (Figure 3(a)) for the perplexity task. Surprisingly, the predictive performance remains remarkably stable even after removing up to 90% of the edges, with minimal degradation. Notably, even under extreme sparsity where only 1% of the original edges are retained, the neural topology still enables effective prediction of perplexity, achieving above 0.6 $R^{2}$ that is still higher than activation probing.
@@
The above experiments suggest that most of the predictive signal resides in a small subset of strong connections, making it possible to significantly reduce the number of edges while preserving nearly all critical topological information. Leveraging this insight, we scale up graph probing to much larger models by operating on sparsified neural topology. While our earlier results focused on models with fewer than 0.5B parameters, we now train probes on sparse graphs derived from LLMs with up to 14B parameters (GPT2-large, Pythia-{1.4B, 2.8B}, Qwen2.5-{3B, 7B, 14B}). As shown in Figure 3(b), graph probing continues to exhibit strong regression performance across all six large models, achieving a maximum $R^{2}$ of over 0.91, providing compelling evidence that the relationship between neural topology and language modeling performance is universal across model sizes. Particularly, the gap between baselines and graph probing remains as substantial as 92.6%, confirming the informativeness of neural topology.
Figure 4: (a-b) Occurrence frequency of hub nodes in (a) Qwen2.5-0.5B and (b) Qwen2.5-1.5B on MMLU benchmark. (c) Accuracy on MMLU benchmark of Qwen2.5 models (0.5B, 1.5B, 3B, 7B, 14B) under different interventions of top 1% neurons.
@@
Probing different layers. We further train probes on neural topology derived from different layers of GPT-2, Pythia-160M, and Qwen2.5-0.5B models. Figure 3(c) reveals that while topologies from all layers are predictive of LLM performance in terms of perplexity, the middle layers are consistently the most informative across all three models, with $R^{2}$ values exceeding 0.8. This finding aligns with prior work identifying the middle layers of LLMs as the hub of semantic processing, whereas the initial and final layers are more specialized for {de, re}-tokenization (Gurnee and Tegmark, 2024; Tuckute et al., 2024).